Blue Water Equipment Company uses the aging approach to estimate bad debt expense at the end of each accounting year. The balance of each account receivable is aged on the basis of the three time periods as follows:

  1. Not yet due, $192,000
  2. Up to one year past due, $12,000
  3. More than one year past due, $3,200

Experience has shown that for each age group, the average loss rate on the amount of the receivable at year-end due to uncollectability is

(1) 1 percent

(2) 3.5 percent, and

(3) 15 percent, respectively

At December 31, 2018 (end of current accounting year), the allowance for doubtful account balance was $1840 (credit).

  1. Determine the amount of estimated uncollectible accounts receivable at the end of 2018
  2. Determine the amount of bad debt expense that should be reported on the income statement
  3. Determine the amount of net accounts receivable that should be reported on the balance sheet at year-end
Answer #1
a
Age of receivables Amount Percent uncollectible Amount uncollectible
Not yet due 192000 1% 1920
Up to one year past due 12000 3.50% 420
More than one year past due 3200 15% 480
Total 207200 2820
Amount of estimated uncollectible accounts receivable = $2820
b
Amount of bad debt expense = 2820-1840= $980
c
Amount of net accounts receivable = 207200-2820= $204380